OpenStudy (anonymous):

IN A RIGHT TRIANGLE, AN ACUTE ANGLE HAS A TANGENT OF 1/3. FIND THE VALUE OF SEC AND SIN OF THIS ANGLE. USE PYTHAGOREAN IDENTITIES I got sec= (radical 10)/3 I can't figure out how to find sin??

OpenStudy (wolf1728):

 By Pythagorean Theorem hypotenuse^2 = opp^2+adj^2 hypotenuse^2 = 9 + 1 hypotenuse^2 = 10 hypotenuse = sqrt 10 sine = opp/hyp sine = 1 / sqrt(10)

OpenStudy (anonymous):

isnt the hypotenuse (radical 10)? because secant is (hypotenuse/adjacent) and my answer was (radical 10) / 3

OpenStudy (wolf1728):

Yes, doesn't radical 10 = square root of 10?

OpenStudy (wolf1728):

secant = hyp/adj secant = sq root (10) / 3 or radical 10 / 3

OpenStudy (anonymous):

ok thank you

OpenStudy (wolf1728):

u r welcome